### v2

# A Bayesian Nonparametric Stochastic Block Model for Directed Acyclic Graphs

Abstract
Random graphs have been widely used in statistics, for example in network and social interaction analysis. In some applications, data may contain an inherent hierarchical ordering among its vertices, which prevents any directed edge between pairs of vertices that do not respect this order. For example, in bibliometrics, older papers cannot cite newer ones. In such situations, the resulting graph forms a Directed Acyclic Graph. In this article, we propose an extension of the popular Stochastic Block Model (SBM) to account for the presence of a latent hierarchical ordering in the data. The proposed approach includes a topological ordering in the likelihood of the model, which allows a directed edge to have positive probability only if the corresponding pair of vertices respect the order. This latent ordering is treated as an unknown parameter and endowed with a prior distribution. We describe how to formalize the model and perform posterior inference for a Bayesian nonparametric version of the SBM in which both the hierarchical ordering and the number of latent blocks are learnt from the data. Finally, an illustration with a real-world dataset from bibliometrics is presented. Additional supplementary materials are available online.
Key Words: Topological Ordering, Directed Acyclic Graphs, Stochastic Block Model, Markov chain Monte Carlo, Pitman-Yor process.
1 Introduction
Random graphs have been widely used in statistics and computer science over the last years to model network, interaction and bibliometric data, just to mention a few among many fields of applications. For example, in social networks data analysis, datasets collecting friendships or exchange of messages among users can be modelled as random graphs, having users as vertices, and edges between two users if they are friends or have exchanged messages. As another example, citation datasets in biliometrics analysis can be modelled as random graphs, having articles as vertices, and directed edges between a pair of articles if the former article cites the latter one.
In some applications, the dataset, and corresponding graph, can contain an inherent hierarchical order among its vertices. This means that directed edges can be present only between vertices in which the former appears before the latter one in this order. As an example, when modelling citation networks among journal articles, (Ji and Jin, 2016), articles published many years ago cannot cite recent articles. As another example, in a network having the employees of a big company as vertices and directed edges from the supervisor to the supervisee, a directed path from one vertex to another can be present only if the former is higher than the latter one in the company hierarchy. Similarly, the same structure appears in a dataset having directed edges from PhD supervisors to their student. In all these applications, the observed graph is a directed graph that cannot display any directed cycles (i.e. any directed path in the graph having the same initial and final vertex). Hence, it is a Directed Acyclic Graph (DAG) and includes an inherent unknown topological ordering among its vertices, i.e. a linear ordering $\prec$ of the vertex set $\mathcal{V}$ such that every pair $(p,q)\in\mathcal{V}\times\mathcal{V}$ can belong to the edge set $E$ only if $p\prec q$.
When analyzing a given network dataset that is a DAG, if the latent topological ordering is not properly accounted for, the statistical models can assign positive probability to ‘impossible’ edges, i.e. directed edges in which the first vertex appears later than the second one in the topological order. If this is the case for many edges, then the resulting estimates from the model can deteriorate dramatically. This is especially true for Bayesian models, and the same phenomenon arises other statistical contexts, for example in presence of structural zeros when modeling tabular data, Manrique-Vallier and Reiter (2014), in which assigning, whatever small, positive probability to ‘impossible’ cells can deteriorate estimates and bias results.
In this work, we propose an approach to extend Bayesian models for random graphs to graphs that are DAGs and contain a latent topological ordering among their vertices. Specifically, we focus on extending the popular Stochastic Block Model (SBM), (Holland et al., 1983), to the case of an observed DAG. The SBM is among the most popular random graphs models for vertex clustering and community detection. This model was initially introduced in (Holland et al., 1983) and (Wang and Wong, 1987) and its statistical formalization as latent models was then described in (Snijders and Nowicki, 1997) and (Nowicki and Snijders, 2001). The model further gained popularity with the incorporation of degree correction (Karrer and Newman, 2011). In online Appendices A.1 and A.2, we provide a brief review of this model, its generalizations and approaches to posterior inference. For recent comprehensive reviews, the reader is also referred to (Abbe, 2018) and (Lee and Wilkinson, 2019), which focus on theoretical results and modeling approaches, respectively. When extending the SBM to model a given DAG, the underlying topological order is unknown, and not necessarily unique. Therefore, it will be included in the likelihood of the model as one of the unknown parameters, endowed with a prior and inferred a posteriori, together with the community memberships and other parameters of the model.
The background literature regarding the use of the SBM for DAGs is quite limited and mainly consists of an unpublished technical report by Lee and Wilkinson (2018), which only presents a preliminary data analysis of a DAG graph using a SBM. Not only was degree correction not incorporated, but there was also no formal consideration of the issue of the number of blocks. Instead, the model was merely fitted with different numbers of blocks separately with no model selection, thus not justifying the interpretations of their results.
The issue of the number of communities or blocks is common with SBMs and clustering in general, and has been reviewed in Table 1 of Lee and Wilkinson (2019). One common approach to this issue is the use of Bayesian nonparametrics, which assigns a prior for the allocation vector and allows the posterior sampler to infer the number of blocks from the data. Kemp et al. (2006) and Schmidt and Mørup (2013) initially considered the infinite relational model, which is a variant of the plain SBM, using a Dirichlet process prior, while Geng et al. (2019) extends this model by the use of a finite regime Pitman-Yor (PY) process (Pitman and Yor, 1997), which is reviewed in the online Appendix A.3. A further generalization is in Legramanti et al. (2022), in which the Gibbs-type prior formulation is used instead of the Pitman-Yor prior, for an application to criminal networks.
In the proposed model, we also use a Bayesian nonparametric version of the SBM, but by adopting the PY process in a different way. This process allows two possible regimes: one in which the total number of blocks is finite (and unknown) and another one in which it increases the sample size. While other works such as Geng et al. (2019) have used similar Bayesian nonparametric formulation for the block membership, they usually focused only on the first of the two regimes. These two regimes imply different asymptotic behaviours of the number of blocks, and the model might be misspecified if only one regime is assumed before fitting to the data. To circumvent this, the choice of regime becomes another part of the proposed model, with a model selection step based on Gibbs variable selection (Carlin and Chib, 1995) embedded in the MCMC sampler.
The proposed approach is illustrated both using simulated data and real datasets from bibliometrics. Specifically, the proposed model including the topological order is compared to a SBM for directed graphs, which does not account for the latent order in the likelihood. The simulation study shows that the latter model assigns positive probability to many ‘impossible’ edges, and brings about substantial biases in the estimates of the overall number of blocks. Then, we present an illustration with two datasets of citation networks, in which it is shown how the model can infer the community structure and latent hierarchical order among vertices.
The rest of this article is organized as follows. The SBM for DAGs is presented in Section 2. Section 3 presents both simulated, Section 3.1, and real data, Section 3.2, illustrations. A final discussion Section 4 concludes the article. The online Appendices contain reviews of the SBM and PY process, a detailed description of the MCMC sampler to perform posterior inference for the SBM for DAG, including both a collapsed and an uncollapsed version of the algorithm, and detailed descriptions of the split and merge proposal and model selection step, other than including some additional information on the likelihood derivation and plots from the real-data illustration.
2 Model
To introduce the model and some notation that will be used in the rest of the paper, let us consider a directed network to be represented as a directed multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the vertex set and $\mathcal{E}$ is the edge set. The size of $\mathcal{V}$, denoted by $n:=|\mathcal{V}|$, is the number of vertices of the graph $\mathcal{G}$. The $n\times{}n$ adjacency matrix of the graph is denoted by $\mathbf{Y}=(Y_{pq})_{1\leq p,q\leq n}$. If there are $y$ directed edges from vertex $p$ to vertex $q$ i.e. $y$ copies of $(p,q)$ in $\mathcal{E}$, then $Y_{pq}=y$, otherwise $Y_{pq}=0$. We also define $\mathbf{Z}_{n}=(Z_{1},\ldots,Z_{n})$ to be the allocation vector of length $n$, where $Z_{p}$ is a label associated to vertex $p$. Two vertices, $p$ and $q$, belong to the same group if and only if $Z_{p}=Z_{q}$. Essentially, $\mathbf{Z}_{n}$ represents the group memberships of the vertices. We assume there are $K_{n}>1$ unique labels displayed in $\mathbf{Z}_{n}$, denoted $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$ hence $K_{n}$ groups of vertices. For notation easiness, now on, we will often write $\mathbb{I}(Z_{p}=i)$, in place of $\mathbb{I}(Z_{p}=Z_{i}^{*})$ and, similarly, given a matrix $\mathbf{C}$ and the event $Z_{p}=Z_{k}^{*}$, $C_{Z_{p}j}$ will denote element $C_{kj}$. Here and in the rest of the paper, $\mathbb{I}(A)$ denotes the indicator function of the event $A$, i.e. $\mathbb{I}(A)=1$ if $A$ is true, 00 otherwise.
The main idea of the classic SBM is that the number of edges $Y_{pq}$ from vertex $p$ to vertex $q$ is independent of that of any other dyad, conditional on their group memberships, $Z_{p}$ and $Z_{q}$, (Holland et al., 1983, Nowicki and Snijders, 2001). Moreover, we will consider the degree-corrected version of the SBM, (Karrer and Newman, 2011), which takes into account the degree heterogeneity of the vertices within the same group, by introducing the vector $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})$ of vertex specific parameters, where $\xi_{p}>0$ is the degree correction factor for vertex $p$. Mathematically, the number of edges between vertex $p$ and $q$ is modeled as
$Y_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\xi}~{}\sim~{}\text{Poisson}(\xi_{p}\xi_{q}C_{Z_{p}Z_{q}}),$ (1)
where $\mathbf{C}=(C_{ij})\in\mathbb{R}_{+}^{K_{n}\times K_{n}}$ is the block matrix. The choice of the Poisson distribution, instead of the Bernoulli distribution, is due to a few reasons, namely, the natural extension to multi-graphs where $\mathbf{Y}$ can take non-negative integer values, the asymptotic equivalence between the edge probability and the expected number of edges for large sparse graphs (Karrer and Newman, 2011), and also the computational simplicity that will be illustrated in our model, especially when considering the degree-corrected version of the model.
In the rest of this work, we will consider a graph $\mathcal{G}$ that is a DAG, and extend the classic SBM to model a DAG. As the graph is directed, $Y_{pq}$ is not necessarily the same as $Y_{qp}$. However, due to the acyclic nature of $\mathcal{G}$, the combinations of the possible values of $(Y_{pq},Y_{qp})$ are restricted to be $(0,0)$, $(0,y)$ or $(y,0)$, for some positive integer $y\in\mathbb{N}$. Also, $Y_{pp}=0$ for all $p=1,2,\ldots,n$, or equivalently the major diagonal contains all zeros. To utilize a unique feature of DAGs, as discussed in Section 1, we define $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{n})$ as the $n$-vector random variable that represents the ordering of $\mathcal{G}$, with the collection of all permutations of $\{1,2,\ldots,n\}$ as the sample space. A value of $\boldsymbol{\sigma}$ is deemed to be topological with respect to $\mathcal{G}$ if it satisfies the definition of topological ordering, i.e. if vertex $p$ topologically precedes vertex $q$, then there cannot be edges from vertex $q$ to vertex $p$. We define several quantities implied from $\boldsymbol{\sigma}$. First, $\boldsymbol{\phi}=(\phi_{1},\ldots,\phi_{n})$ is the “inverse” of $\boldsymbol{\sigma}$, which means that if vertex $p$ is the $r$-th vertex in the topological ordering, we have $\sigma_{r}=p$ and $\phi_{p}=r$. Essentially, $\boldsymbol{\phi}$ contains the position of each vertex in $\boldsymbol{\sigma}$. For convenience, without confusion, we write $\boldsymbol{\sigma}^{-1}=\boldsymbol{\phi}$ and $\boldsymbol{\phi}^{-1}=\boldsymbol{\sigma}$. Second, $\mathbf{Z}_{n}^{\boldsymbol{\sigma}}=\left(Z^{\boldsymbol{\sigma}}_{1},Z^{\boldsymbol{\sigma}}_{2},\ldots,Z^{\boldsymbol{\sigma}}_{n}\right)$ is the reordered allocation vector, where $Z^{\boldsymbol{\sigma}}_{p}=Z_{\sigma_{p}}$, where $Z_{\sigma_{p}}$ comes from the allocation vector $\mathbf{Z}_{n}$. In the same way, if we define $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{n})$ as the vector of degree correction parameters, $\boldsymbol{\xi}^{\boldsymbol{\sigma}}=\left(\xi^{\boldsymbol{\sigma}}_{1},\xi^{\boldsymbol{\sigma}}_{2},\ldots,\xi^{\boldsymbol{\sigma}}_{n}\right)$ is its reordered version. Lastly, $\mathbf{Y}^{\boldsymbol{\sigma}}$ is the adjacency matrix reordered by $\boldsymbol{\sigma}$ for the columns and rows of $\mathbf{Y}$ simultaneously, such that $Y_{pq}^{\boldsymbol{\sigma}}=Y_{\sigma_{p}\sigma_{q}}$.
The central component of the SBM is the distribution assumption about each dyad of $\mathcal{G}$, essentially each element of $\mathbf{Y}$ or equivalently $\mathbf{Y}^{\boldsymbol{\sigma}}$. As $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular for a $\boldsymbol{\sigma}$ that is topological, we require that, for all dyads $(p,q)$ where $1\leq p<q\leq n$,
$Y^{\boldsymbol{\sigma}}_{qp}=0,\quad\quad Y^{\boldsymbol{\sigma}}_{pq}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi}$ $\sim\text{Poisson}\left(\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}C_{Z^{\boldsymbol{\sigma}}_{p}Z^{\boldsymbol{\sigma}}_{q}}\right),$ (2)
where $\mathbf{C}=(C_{ij})\in\mathbb{R}_{+}^{K_{n}\times K_{n}}$ is the block matrix.
As $\mathbf{Z}_{n}$ is unknown prior to fitting the model, it will be treated as a vector of latent variables and assigned a prior. Specifically, we assume that $\mathbf{Z}_{n}$ are the first of $n$ elements of an exchangeable sequence driven by a PY process prior. The parametrization of this prior will be detailed in Section 2.2. Given this prior choice, the vector $\mathbf{Z}_{n}$ will display $K_{n}$ distinct values $(Z_{1}^{*},\ldots,Z_{K_{n}}^{*})$, which appear in $\mathbf{Z}_{n}$ with frequencies $\mathbf{N}=(N_{1},\ldots,N_{K_{n}})$. The $i$-th element $N_{i}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=Z_{i}^{*})$ is the number of vertices in group $i$, displaying label $Z_{i}^{*}$. As previously mentioned, we write $\mathbb{I}(Z_{p}=i)$, in place of $\mathbb{I}(Z_{p}=Z_{i}^{*})$. Also, we derive the $K_{n}\times K_{n}$ edge matrix between groups from $\mathbf{Z}_{n}$ and $\mathbf{Y}$, denoted by $\mathbf{E}$, where $E_{ij}=\sum_{p=1}^{n}\sum_{q=1}^{n}Y_{pq}\mathbb{I}(Z_{p}=i,Z_{q}=j)$.
Next, we define a $K_{n}\times{}K_{n}$ matrix $\mathbf{M}$, where $M_{ij}$ is the number of dyads $(p,q)$ between groups $i$ and $j$ such that vertex $p$ is topologically in front of vertex $q$. Mathematically,
$M_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z_{\sigma_{p}}=i,Z_{\sigma_{q}}=j\right)=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}\left(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j\right).$
2.1 Likelihood
With all required quantities defined, we will derive the likelihood with observed $\mathbf{Y}$ given $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$. We first check that $\boldsymbol{\sigma}$ is topological, or equivalently $\mathbf{Y}^{\boldsymbol{\sigma}}$ is upper triangular i.e. $Y^{\boldsymbol{\sigma}}_{qp}=0$ for all dyads $(p,q)$ where $1\leq p<q\leq n$, otherwise the likelihood is 0. Once $\boldsymbol{\sigma}$ is checked to be topological, using Equation 2, the observed data likelihood is
$\mathbb{P}(\mathbf{Y}|\mathbf{C},\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi})=\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})\times\bar{Y}\times\prod_{p=1}^{n}\prod_{q=1}^{n}\left(\xi_{p}\xi_{q}\right)^{Y_{pq}}\times\prod_{i=1}^{K_{n}}\prod_{j=1}^{K_{n}}e^{-C_{ij}M_{ij}}C_{ij}^{E_{ij}},$ (3)
where $\bar{Y}=\prod_{p=1}^{n-1}\prod_{q=p+1}^{n}\left(Y^{\boldsymbol{\sigma}}_{pq}!\right)^{-1}$, ${}\displaystyle M_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}\xi^{\boldsymbol{\sigma}}_{p}\xi^{\boldsymbol{\sigma}}_{q}\mathbb{I}(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j)$, and ${}\displaystyle E_{ij}=\sum_{p=1}^{n-1}\sum_{q=p+1}^{n}Y^{\boldsymbol{\sigma}}_{pq}\mathbb{I}(Z^{\boldsymbol{\sigma}}_{p}=i,Z^{\boldsymbol{\sigma}}_{q}=j)$. Detailed derivations are in the online Appendix B. Likelihood 3 is influenced by $\mathbf{Z}_{n}$ and $\boldsymbol{\sigma}$ through the two matrices $\mathbf{E}$ and $\mathbf{M}$.
2.2 Priors and posterior density
We shall assign independent priors one by one to $\mathbf{C}$, $\mathbf{Z}_{n}$, $\boldsymbol{\sigma}$ and $\boldsymbol{\xi}$, in order to carry out inference within the Bayesian framework. In the subsequent calculations, some additional parameters of the priors used will be included in the notation.
For $\mathbf{C}$, we assume each $C_{ij}$ is a priori independent and identically distributed according to the Gamma$(a,b)$ distribution, where $a$ and $b$ are the positive shape and rate parameters, respectively. This enables $\mathbf{C}$ to be integrated out, to obtain
[…]
$\quad=\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})\times\bar{Y}\times\prod_{p=1}^{n}\prod_{q=1}^{n}\left(\xi_{p}\xi_{q}\right)^{Y_{pq}}\times\left(\frac{b^{a}}{\Gamma(a)}\right)^{K_{n}^{2}}\prod_{i=1}^{K_{n}}\prod_{j=1}^{K_{n}}\frac{\Gamma(E_{ij}+a)}{(M_{ij}+b)^{E_{ij}+a}}$ (4)
Independent and relatively uninformative gamma prior distributions are assigned to the parameters $a$ and $b$, as well as the components of $\boldsymbol{\xi}$.
For $\boldsymbol{\sigma}$, we assign a uniform prior to all permutations of $\{1,2,\ldots,n\}$ i.e. $\pi(\boldsymbol{\sigma})=(n!)^{-1}$. There is no issue with an ordering that is not topological having a positive prior probability, as such an ordering will result in $\mathbb{I}(\mathbf{Y}^{\boldsymbol{\sigma}}\text{ upper tri.})$. ) and the likelihood (4) being equal to 0.
For $\mathbf{Z}_{n}$, we assume that it comes from an exchangeable sequence $(Z_{p})_{p\in\mathbb{N}}$ driven by a Pitman-Yor process, Pitman and Yor (1997), i.e. let $Z_{p}|P\stackrel{{\scriptstyle iid}}{{\sim}}P$ for all $p\in\mathbb{N}$ and $P\sim\text{PY}(\alpha,\theta,P_{0})$. The Pitman-Yor process, reviewed in the online Appendix A.3, is a distribution for an unknown probability distribution $P$ and is parametrized by three hyperparameters $(\alpha,\theta,P_{0})$, where $P_{0}$, called base distribution, is a distribution on the sample space, and $\alpha$ and $\theta$ are two scalars satisfying either: 1) $10\leq\alpha<$ 1 and $\theta\geq-\alpha$; 2) $\alpha<0$ and $\theta=k|\alpha|$ for $k\in\mathbb{N}$. Case 1) corresponds to the infinite regime, in which $P$ has an infinite number of support points, and the sequence $(Z_{p})_{p\in\mathbb{N}}$ will display infinite number of distinct values. Case 2) is the finite regime and $P$ has $k$ support points, which is also the total number of distinct values in $(Z_{p})_{p\in\mathbb{N}}$. The Dirichlet process corresponds to the special case $\alpha=0$. When considering the finite regime, i.e. when $\alpha<0$, we apply the re-parametrization $(\alpha,\theta)\to(\gamma,k)$, with $\gamma:=|\alpha|>0$ and $k:=\theta/|\alpha|\in\mathbb{N}$, and assign a prior to $(\gamma,k)\in\mathbb{R}_{+}\times\mathbb{N}$.
As detailed in the online Appendix A.3, it is possible to compute the marginal likelihood (when $P$ is marginalized out) of a sample $\mathbf{Z}_{n}$ driven by the PY process, which we denoted by $\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})$, where $\boldsymbol{\eta}_{r}$ is a parameter vector of length 2, dependent on the choice of regime $r\in\{0,1\}$, either infinite $\boldsymbol{\eta}_{0}=(\alpha,\theta)$ or finite $\boldsymbol{\eta}_{1}=(\gamma,k)$. This in turn requires the specification of the prior of $\boldsymbol{\eta}_{r}$ under both regimes. Under the infinite regime, $r=0$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{0}=(\alpha,\theta)$, and we assume that $\alpha\sim\text{Uniform}[0,1]$ and $\theta+\alpha$ follows a Gamma distribution. This choice of dependent priors for $\alpha$ and $\theta$ is done in order to include all possible values in parameter space. Indeed, when $10<\alpha<$ 1, the range of possible values of $\theta$ is $\theta>-\alpha$. Under the finite regime, $r=1$ and $\boldsymbol{\eta}_{r}=\boldsymbol{\eta}_{1}=(\gamma,k)$, and we assume that $\gamma$ and $k$ are independent a priori, $\gamma$ follows a Gamma distribution, and $k$ follows a truncated negative binomial distribution, with parameters $a_{k}$ and $b_{k}$, and density
$\mathbb{P}(k=k^{\prime}|a_{k},b_{k})=\left(1-b_{k}^{a_{k}}\right)^{-1}\times\frac{\Gamma(k^{\prime}+a_{k})}{\Gamma(a_{k})k^{\prime}!}~{}b_{k}^{a_{k}}(1-b_{k})^{k^{\prime}},\qquad k^{\prime}=1,2,\ldots$
where $\Gamma(\cdot)$ is the gamma function. The factor $(1-b_{k}^{a_{k}})^{-1}$ is due to the truncation of 00 from the original support of the negative binomial distribution. The negative binomial is over-dispersed, allowing possibly high variance, hence high prior uncertainty of the value of $k$.
The regime-dependent parameters and their priors are introduced this way because, ultimately, we want to enable model selection of the regime, which in turn requires the prior of $r$, denoted by $\mathbb{P}(r)$. The boundary cases $\mathbb{P}(r=0)=1$ and $\mathbb{P}(r=1)=1$ represent staying within the infinite and finite regimes, respectively, while the incorporation of the model selection in the MCMC sampler allows the data to select the more suitable among the two regimes. The model selection step, implemented within the MCMC sampler, follows the algorithm of Carlin and Chib (1995). Moreover, the prior choice for $r$ allows unbalanced weights between models. This might improve mixing of the model selection step, as shown in Friel and Pettitt (2008), in the context of reversible jump MCMC. At the end of Section 3.1, in Table 1, we show results from a simulated study in which the posterior sampler seems to be able to recover the correct regime between the finite and infinite ones. A detailed description of the model selection step can be found in the online Appendix C.2.
As all priors required have been specified, the joint posterior of $\mathbf{Z}_{n}$, $\boldsymbol{\sigma}$, $\boldsymbol{\eta}_{r}$, $a$ and $b$ (and $r$), up to a proportionality constant, is readily available:
$\mathbb{P}(\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},\boldsymbol{\eta}_{r},a,b,r|Y)\propto\mathbb{P}(\mathbf{Y}|\mathbf{Z}_{n},\boldsymbol{\sigma},\boldsymbol{\xi},a,b)\mathbb{P}(\mathbf{Z}_{n}|\boldsymbol{\eta}_{r})\mathbb{P}(\boldsymbol{\sigma})\mathbb{P}(\boldsymbol{\xi})\mathbb{P}(\boldsymbol{\eta}_{r}|r)\mathbb{P}(a)\mathbb{P}(b)\mathbb{P}(r).$ (5)
We carry out Bayesian inference by sampling from this joint posterior using MCMC. The full detail description of the MCMC sampler is available in the online Appendix C.1.
3 Application
3.1 Simulated data
This section presents an illustrative simulation study in which the proposed DAG-SBM (Equation 2), is compared with the SBM for a directed graph (Equation 1, referred to as the directed SBM). The purpose of this study is to show how fitting the directed SBM without considering the inherent hierarchical ordering among vertices can result in assigning positive posterior probability to many “impossible” edges, and substantial bias in the estimate of the overall number of blocks in the model.
Upon sampling the parameters from their priors specified in Section 2.2, adjacency matrices are sampled from the DAG-SBM for different sample sizes $n=\{250$,$500$,1000}1000\}1000 } and combinations of the Pitman-Yor parameters, $(\alpha,\theta)=\{(-1,10)$, $(0,1)$, $(0.1,1)$, (0.6,1)}(0.6,1)\}( 0.6 , 1 ) }, thus including both finite and infinite regimes. Both the DAG-SBM and directed SBM are fitted to the resulting adjacency matrices, also under both the finite and infinite regimes.
Figure 1: Actual adjacency matrix (left) for data simulated from the DAG-SBM (Equation 2) with $n=250$ and different values of $(\alpha,\theta)$, and estimated probabilities by the directed SBM (Equation 1) under the finite (middle column) and infinite (right) regime, ordered by true $\boldsymbol{\sigma}$. The middle and right plots correspond to the results in finite and infinite regimes in the corresponding bottom left panels of Figure 2. The dashed line is the major diagonal.
Figure 1 plots the heat maps of the actual adjacency matrix $\mathbf{Y}^{\boldsymbol{\sigma}}$ with $n=250$ (not shown for $n=500$ and $n=1000$), ordered according to the true topological ordering, alongside with the posterior point estimates of $\mathbf{Y}^{\boldsymbol{\sigma}}$, for each simulated data set fit by the directed SBM. The true adjacency matrices in the first column are necessarily upper triangular, given that they have been reordered according to the true $\boldsymbol{\sigma}$. However, from the second and third columns, it is evident that many edges in the lower triangular sections have positive posterior probability, while it should be equal to zero. This is especially true for vertices having either high in-degree or out-degree counts. This is due to the fact that the directed SBM is row and column exchangeable and vertices with high in-degree tends to have also high out-degree counts. However, when a hierarchical ordering is present, the model is not exchangeable anymore and the out- and in-degrees of each vertex should be negatively correlated random variables, with vertices at the beginning the topological order having higher out-degree than in-degree counts, while the opposite for vertices at the end of the topological order.
Figures 2 displays the posterior histograms of the number of blocks $K_{n}$ estimated using the DAG-SBM and the directed SBM, for the same sets of simulated data. On the one hand, the directed SBM, which does not include the topological ordering in the likelihood, has the tendency of heavily underestimating or overestimating $K_{n}$. The increase in the number of blocks also negatively affects the computational costs of the algorithm, compared to the DAG version. A comparison of the computational times between the two models is provided in the online Appendix D.1. On the other hand, the DAG-SBM, which accounts for the topological ordering, provides reasonable estimates of $K_{n}$ across different sample sizes and for both regimes of the Pitman-Yor parameters. Furthermore, for the DAG-SBM, the model selection step in the MCMC sampler successfully recovers the true regime of the Pitman-Yor parameters. Indeed, Table 1 displays the estimated Bayes factor for the model selection step between the finite and infinite regime, for a range of different values of the hyperparameters of the Pitman-Yor process. The estimated Bayes factor suggests that the posterior sampler is capable at recovering the correct regime from the data, particularly when $\alpha$ is far from 00, and being generally indifferent between the regimes when $\alpha$ is close to 00.
Figure 2: Posterior histogram of $K_{n}$ for data simulated from the DAG-SBM (Equation 2), with different values of $(\alpha,\theta,n)$, fitted by both the DAG-SBM (top row) and directed SBM (Equation 1, bottom row), under the finite (solid) and infinite (stripe) regimes. The vertical dashed line is the true value of $K_{n}$.
$\alpha$ $\theta$ $n=250$ $n=500$ $n=1000$
$-1$ $10$ $8.30$ $612.58$ $2.72$
00 $1$ $2.56$ $0.85$ $19.12$
$0.1$ $1$ $1.11$ $0.40$ $0.89$
$0.6$ $1$ $7.93\times 10^{-4}$ $7.62\times 10^{-7}$ $1.92\times 10^{-11}$
Table 1: Estimated Bayes factor for the finite regime to the infinite regime when applying the model selection in the MCMC sampler for the DAG-SBM to the simulated data, for different values of $(\alpha,\theta,n)$.
Finally, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is sufficiently good that the trace plots are not shown. The posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 3, with the rows in the true topological ordering. The alignment suggests that this latent ordering can be well recovered without identifiability issues.
Figure 3: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for $n=250$ and different values of $(\alpha,\theta)$ for data simulated from the DAG-SBM (Equation 2), fitted by the DAG-SBM under the true regime according to $\alpha$. Each row of colours is the posterior density of a component of $\boldsymbol{\phi}$.
3.2 SNA citation network
We first look at the citation network analysed by Lee and Wilkinson (2018). It contains 1118 citations (edges) between 135 articles (vertices) which are related to social network analysis (SNA). We shall call this data the SNA citation network hereafter.
The MCMC sampler for the DAG-SBM outlined in the online Appendix C was applied, with 20000 iterations obtained after a burn-in period of 100000 and thinning of 2000 i.e. keeping 1 sample every 2000 iterations, with $a\sim\text{Gamma}(1,0.01)$, $b\sim\text{Gamma}(1,0.01)$. Such settings were used three times, once assuming the infinite regime $(r=0)$ with $\theta+\alpha\sim\text{Gamma}(1,0.01)$, once assuming the finite regime $(r=1)$ with $\gamma\sim\text{Gamma}(1,0.01)$ and $k\sim$ truncated negative binomial$(1,0.01)$, and once for model selection with $\mathbb{P}(r=1)=0.2$ i.e. $\mathbb{P}(r=0)=0.8$. All three runs were performed on a Linux machine with Intel Core i7-7700 Processor (3.6GHz), and took 0.0089, 0.0077, and 0.0083 seconds per iteration, respectively.
Figures 4 to 9 show some key inference results, except for the trace plots of the parameters, which are in the online Appendix D.2. Of more importance is the posterior density (or mass function in the case of $K_{n}$ and $k$) in Figures 4 and 5. On one hand, in the panels for $k$ and $\gamma$, the infinite regime is naturally missing, while the results by the finite regime and model selection coincide as expected. Similarly, in the panels for $\theta$ and $\alpha$, the finite regime is naturally missing, while the results by the infinite regime and model selection coincide as expected. On the other hand, the panels for $K_{n}$, $a$ and $b$ illustrate that the posterior density from model selection is essentially a weighted average of that of the two regimes.
The departure of the posterior densities of $\alpha$ and $\gamma$ from 0 suggests that either the infinite regime or the finite regime is preferred to their shared boundary i.e. the Dirichlet process when $\alpha=\gamma=0$. Between the two regimes, with $\mathbb{P}(r=1)=0.2$ resulting in $\mathbb{P}(r=1|\mathbf{Y})=0.5474$, the Bayes factor $B_{10}=\displaystyle\frac{\mathbb{P}(r=1|\mathbf{Y})}{\mathbb{P}(r=0|\mathbf{Y})}\left/\frac{\mathbb{P}(r=1)}{\mathbb{P}(r=0)}\right.=4.8368$, suggesting a slight preference for the finite regime, for this citation network.
Figure 4: Posterior histogram of $K_{n}$ and $k$, for the finite (solid) and infinite (stripe) regimes of the DAG-SBM, and model selection (crosshatch), for the SNA citation network.
Figure 5: Posterior density of the parameters for the infinite (dotted) and finite (solid) regime of the DAG-SBM, and model selection (dashed), for the SNA citation network.
Figure 6: Joint posterior of $\log\gamma$ and $\log k$ for the finite regime of the DAG-SBM, for the SNA citation network.
The high number of thinning is mainly due to the Markov chain for the finite regime occasionally getting stuck at $K_{n}=4$, which is in turn due to the mixing of $k$ and $\gamma$, and the skewness of their joint posterior, which is plotted on log scale in Figure 6. There is a dense cloud of points towards the bottom right of the plot, indicating high posterior density around small values of $k$ (and high values of $\gamma$), thus limiting $K_{n}$ to grow. However, such issue of getting stuck at small values of $k$ and $K_{n}$ does not exist for either the infinite regime or the application to the larger data set in the next subsection.
Next, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is sufficiently good and the posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 7, with the rows themselves in an arbitrary topological ordering. As the citations i.e. the edges in the DAG in general go from more recent works to older ones, the top (bottom) rows, which correspond to topologically earlier (later) vertices, are indeed the more recent (older) articles in the network. The coloured part of each row represents the support of the posterior of the corresponding article, and in general is narrower towards the bottom. An interpretation is that chronologically older works are likely to cite between themselves and get cited by later ones, thus limiting their positions in $\boldsymbol{\sigma}$.
Figure 7: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network. Each row of colours is the posterior density of a component of $\boldsymbol{\phi}$. The black dots are the mean positions.
Figure 8: The co-clustering matrix (spectrum) and the adjacency matrix (dots) for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network. The vertices are clustered (dashed lines) according to the point estimate $\hat{\mathbf{Z}}_{n}$.
Figure 9: Network diagram according to the point estimate of the allocation vector, $\hat{\mathbf{Z}}_{n}$, for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network.
To provide a posterior point estimate of $\mathbf{Z}_{n}$, we follow the clustering approach introduced in (Meilă, 2007) and further discussed in (Wade and Ghahramani, 2018). Specifically, the point estimate, denoted by $\hat{\mathbf{Z}}_{n}$, is obtained using a decision theoretic approach, by minimising with respect to the posterior distribution a loss function on the space of allocation vectors,
$\hat{\mathbf{Z}}_{n}=\underset{{\tilde{\mathbf{Z}}_{n}}}{\text{argmin}}~{}\mathbb{E}\left[L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})|\mathbf{Y}\right]=\underset{{\tilde{\mathbf{Z}}_{n}}}{\text{argmin}}\sum_{\mathbf{Z}_{n}}L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})\mathbb{P}(\mathbf{Z}_{n}|\mathbf{Y}).$ (6)
For the loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$, (Meilă, 2007) chose the Variation of Information (VI), defined as
$\text{VI}(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})=\sum_{i=1}^{K_{n}}\frac{n_{i+}}{n}\log\left(\frac{n_{i+}}{n}\right)+\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{+j}}{n}\log\left(\frac{n_{+j}}{n}\right)-2\sum_{i=1}^{K_{n}}\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{ij}}{n}\log\left(\frac{n_{ij}}{n}\right)$
where the counts variables are defined as $n_{ij}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=i,\tilde{Z}_{p}=j)$, $n_{i+}=\sum_{j=1}^{\tilde{K}_{n}}n_{ij}$, and $n_{+j}=\sum_{j=1}^{K_{n}}n_{ij}$. The loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$ can be seen as a distance between $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$, which can be computed even if $K_{n}\neq\tilde{K}_{n}$ i.e. the numbers of groups implied by $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$ are different. Using the SALSO algorithm (Dahl et al., 2022) to minimise the loss function and obtaining the point estimate $\hat{\mathbf{Z}}_{n}$ for the SNA citation network, we plot the co-clustering and adjacency matrices in Figure 8, and the network diagram in Figure 9.
3.3 Statistics citation network
We apply the model to a second citation network. The original data analysed by (Ji and Jin, 2016) contains 5722 citations between 3248 articles in the top statistics journals. We only consider the largest connected component, and remove one of the two edges in every pair of cyclic citations (there were only 9 pairs of these). Upon the data cleaning, we arrive at a citation network that is a DAG and contains 5563 citations between 2248 articles. We refer to this data as the statistics citation network hereafter.
The MCMC sampler was applied with $1.5\times 10^{4}$ iterations obtained after a burn-in period of $5.5\times 10^{4}$ and thinning of 10, with the same priors for $a$, $b$, $\alpha$, $\theta$, $k$ and $\gamma$ as those in Section 3.2. For model selection, even with $\mathbb{P}(r=0)=0.99$, the whole chain stays in the finite regime i.e. $\mathbb{P}(r=1|\mathbf{Y})=1$. Therefore, we shall report the results under the finite regime only, as it is heavily preferred to the infinite counterpart for this network. The trace plots and posterior densities are provided in the online Appendix D.3. The posterior of $K_{n}$ ranges from 92 to 124, which seems reasonable with around 30 articles per cluster.
Similar to Figure 8, the adjacency matrix for the statistics citation network is plotted in Figure 10, with the vertices clustered according to the point estimate. The co-clustering matrix is not being overlaid here to preserve the figure readability. The concentration of the black dots along the major block diagonal suggests that most groups are closely knitted. On the other hand, there are some concentrated blocks which are off-diagonal and asymmetric, indicating high number of one-way citations from one group to another.
Figure 10: The adjacency matrix (black dots) for the finite regime, for the statistics citation network. The vertices are clustered (dashed lines) according to the point estimate $\hat{\mathbf{Z}}_{n}$.
Table 2: Titles of top articles according to topological ordering for selected groups
Group Article
18 “False Discovery” Genovese and Wasserman (AoS, 2004)
18 “False Discovery” Storey (AoS, 2003)
18 “False Discovery” Johnstone (AoS, 2008)
18 “False Discovery” Black (JRSSB, 2004)
18 “False Discovery” Cox and Wong (JRSSB, 2004)
45 “Design and Randomization” Lockwood, Schervish, Gurian and Small (JASA, 2004)
45 “Design and Randomization” Montanari and Ranalli (JASA, 2005)
45 “Design and Randomization” Rubin (JASA, 2005)
45 “Design and Randomization” Barnard, Frangakis, Hill and Rubin (JASA, 2003)
45 “Design and Randomization” Greevy, Silber, Cnaan and Rosenbaum (JASA, 2004)
2 “Wavelet techniques” Averkamp and Houdré (AoS, 2005)
2 “Wavelet techniques” Morris, Vannucci, Brown and Carroll (JASA, 2003)
2 “Wavelet techniques” Yu and Jones (JASA, 2004)
2 “Wavelet techniques” Foster and Stine (JASA, 2004)
2 “Wavelet techniques” Signorini and Jone (JASA, 2004)
71 “Missing Data” Rao, Mingo, Speicher and Edelman (AoS, 2008)
71 “Missing Data” Li, Aragon, Shedden and Agnan (JASA, 2003)
71 “Missing Data” Stute, Xue and Zhu (JASA, 2007)
71 “Missing Data” Liang, Wang, Robins and Carroll (JASA, 2004)
71 “Missing Data” Chen, Leung and Qin (JASA, 2003)
25 “Functional Regression” Crambes, Kneip and Sarda (AoS, 2009)
25 “Functional Regression” Hall and Horowitz (AoS, 2007)
25 “Functional Regression” Cai and Hall (AoS, 2006)
25 “Functional Regression” Wu and Liu (JASA, 2007)
25 “Functional Regression” Mammen and Nielsen (Biometrika, 2003)
65 “Spatial Processes” Wang and Carey (JASA, 2004)
65 “Spatial Processes” Schmidt and O’Hagan (JRSSB, 2003)
65 “Spatial Processes” Schlather, Ribeiro Jr and Diggle (JRSSB, 2004)
65 “Spatial Processes” McElroy and Politis (AoS, 2007)
65 “Spatial Processes” Gelfand, Kim, Sirmans and Banerjee (JASA, 2003)
Table 3: Titles of top 20 articles according to topological ordering
Article
Johnstone (AoS, 2008)
Genovese and Wasserman (AoS, 2005)
Zhang, Siegmund, Ji and Li (Biometrika, 2010)
Averkamp and Houdré (AoS, 2005)
Ishwaran and Rao (JASA, 2003)
Zeng (AoS, 2004)
Beskos, Papaspiliopoulos and Roberts (AoS, 2009)
Schlather and Tawn (Biometrika, 2003)
Mammen and Nielsen (Biometrika, 2003)
Wang and Carey (Biometrika, 2003)
Zhang (JASA, 2003)
Huang (AoS, 2004)
Hu and Rosenberger (JASA, 2003)
Ramsay, Hooker, Campbell and Cao (JRSSB, 2007)
Storey (AoS, 2003)
Lee and Pun (JASA, 2006)
Zhu, Miao and Peng (JASA, 2006)
Barnard, Frangakis, Hill and Rubin (JASA, 2003)
Foster and Stine (JASA, 2004)
Lue (Biometrika, 2004)
Also, we report several groups of top articles according to median topological ordering for some large groups in Table 2. Specifically, for a selection of blocks, we report the top 5 articles in topological order within the block. From the articles’ titles, groups seem to roughly correspond to various topics in statistics. For example, group 18 seems to cover false discovery rates, group 45 design and randomisation, group 2 wavelets, group 71 missing/perturbed data, group 25 function linear regression, and group 65 spatial processes. However, note that the reported group memberships are just point estimates obtained with the MCMC draws, and therefore there might be some possibly large posterior uncertainty in the topological ordering for some of these articles.
Lastly, we report in Table 3 some top articles according the median topological ordering, this time without the group memberships. In general, these articles have high number of citations. However, we also notice that some of these articles might not have a very high number of citations. This is due to the fact that these are articles are cited by some other influential articles towards the end of the topological ordering.
4 Discussion
In this article, we proposed a Bayesian nonparametric SBM for DAGs. Specifically, by conditioning on a latent topological ordering, the likelihood of the data (which is composed of directed edges) becomes that of an undirected graph, i.e. that of an upper diagonal adjecency matrix. The topological ordering is treated as an unknown parameter, endowed with a prior and inferred a posteriori within the MCMC sampler, using a modified Leap-and-Shift proposal. Also, the use of the PY process prior for the allocation vector $\mathbf{Z}_{n}$ allows the model to infer the number of groups $K_{n}$ from the data. Moreover, a model selection step for the two regimes of the PY process can be included within the MCMC sampler.
The model can be generalized in different ways. For example, it can be extended by introducing covariate information, such as the keywords of each document or its publication year. This could be achieved by modelling the degree correction factors with a covariate dependent distribution. Also, in terms of parametrisation, the two regimes of the PY process could be unified so that $\gamma$ and $\alpha$ become one parameter that can take a value between $\infty$- ∞ and $1$. Its posterior density will directly imply which regime is preferred. The main obstacle to overcome here would be the sampling from the non-standard joint parameter space of $\theta$ and $\alpha$ across the two regimes.
Another issue to be resolved is the inference of $k$, which is naturally highly correlated with $K_{n}$, under the finite regime. This is apparent in the parameter trace plots (in online Appendix D.3) for the statistics citation network, while for the SNA citation network the model selection improves the mixing of $k$ in the MCMC. Ideally $k$ is integrated out, but the computations required mean that this is feasible only under certain special cases. Such issue with $k$ remains to be resolved.
There are potential extensions regarding inference procedure and results. Similar to how $\hat{\mathbf{Z}}_{n}$ is computed for $\mathbf{Z}_{n}$, a point estimate could be provided for $\boldsymbol{\sigma}$, but the distance function for the ordering has to be carefully considered. Relatedly, a Mallow’s model prior (Vitelli et al., 2018) could be used for $\boldsymbol{\sigma}$, as opposed to the uniform prior used here, to potentially provide more information to facilitate the inference. Lastly, the derivation of an efficient Variational Bayes algorithm (Blei et al., 2017) would possibly allow the applicability of the proposed model to much larger datasets.
Supplementary Materials
A single archive, available online, contains all of the following supplementary materials:
Appendices:
Background materials of SBM and Pitman-Yor process, likelihood derivations, MCMC algorithms, and additional plots (appendices.pdf).
R package for simulations & MCMC sampler:
R-package “dagsbm” containing code to run the MCMC sampler for the DAG-SBM, and to simulate graphs from the model. The package also contains the dataset of SNA citation network presented in the article. (dagsbm_0.0.7.tar.gz, GNU zipped tar file)
Code:
Script for running the MCMC sampler using the R package. (sim_dc.R)
Results:
Pre-generated results from running sim_dc.R, for both the README and reproducibility checks. (results/*sim_dc_alpha=0.00_theta=1.00_n=250*.rds)
README:
Instructions for reproducing the results in the paper using the R package and code. (README.Rmd)